OpenStudy (anonymous):

A liquid vessel which is initially empty is in the form of an inverted regular hexagonal pyramid of altitude 25 ft. and base edge 10 ft. How much will the surface rise when 6,779 liters of water is added?

OpenStudy (cwrw238):

The volume of a pyramid = (1/3) * altitude* area of the base area of a regular hexagon = (3 sqrt3 a^2) / 2 where a = length of the edgs

OpenStudy (cwrw238):

in your case a = 10, the volume = 6779 and you need to find the altitude L

OpenStudy (cwrw238):

so we have 6779 = (1/3) * L * (3 sqrt3 10^2) / 2 solve for L

OpenStudy (anonymous):

3 sqrt a^2 is only applicable to what regular polygons? my professor teaches a different formula

OpenStudy (cwrw238):

(3 sqrt3 * a^2 )/ 2 is applicable to a regular hexagon

OpenStudy (anonymous):

oh, I get L - 78.28

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

thats what i get too

OpenStudy (cwrw238):

Oh - but the vessel is only 25 ft high!!

OpenStudy (anonymous):

lol it says 12 ft as the answer at the back of the book

OpenStudy (cwrw238):

theres something wrong somewhere....

OpenStudy (anonymous):

well you need to find the height of the smaller pyramid you make by reducing the volume of the larger pyramid by 6779 I think

OpenStudy (anonymous):

I think you need to convert 6779 liters to cubic ft

OpenStudy (cwrw238):

oh i see whats wrong - the vessel is inverted - so as you say you have to find the volume of smaller pyramid

OpenStudy (cwrw238):

sorry guy - i messed up there.

OpenStudy (anonymous):

no you need to get the height difference i think

OpenStudy (cwrw238):

yeah - its 6779 liters not cubic feet

OpenStudy (anonymous):

oh yeah you're right the volume lol

OpenStudy (cwrw238):



OpenStudy (cwrw238):

we need conversion factor for liters to cu ft

OpenStudy (anonymous):

that's 0.0353ft^3/L

OpenStudy (anonymous):

it's 239.3 cubic ft

OpenStudy (cwrw238):

ok

OpenStudy (cwrw238):